OpenStudy (anonymous):

ok last one plz help 7x-y=8 9x+y=24 using combination

OpenStudy (anonymous):

since you already have opposite signs for y terms...they will cancel each other out when you add them

OpenStudy (anonymous):

yeah i got that

OpenStudy (anonymous):

7x - y =8 9x + y = 24 __________________ 16x = 32 ___ __ 16 16 x= 2

OpenStudy (anonymous):

now plug 2 in for x in one of the original equations to get y

OpenStudy (anonymous):

7(2)-y=8 14-y=8 add y then subtract 8 or subtract 14 and divide by -1 on both sides -y=-6 y = 6

OpenStudy (anonymous):

the solution is (2,6)

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

welcome